OpenStudy (anonymous):

A wheel is rotating at 5 radians/sec, and the wheel has a 62-inch diameter. To the nearest foot per minute, what is the speed of a point on the rim?

OpenStudy (yttrium):

It's linear velocity, right?

OpenStudy (yttrium):

V = rw where V is the linear velocity r is the radius which is 62/2 w is the angular velocity which is 5rads/s

OpenStudy (yttrium):

Hence you will arrive at \[V = 31(5)\] \[V = 155\] but it is in terms inch per sec. So we'll have to convert it

OpenStudy (yttrium):

\[V = 155 (\frac{ 1ft }{ 12inch })(\frac{ 60s }{ 1\min })\] Hence we wil have V = 775ft/min

OpenStudy (anonymous):

Thank you so much!